A power plant delivers 900.0 MW of usable energy and dumps 2100.0 MW of waste heat into a nearby river at 15*C. What is the minimum temperature that the combustion locaiton could possibly have?
Answer #1

Here ,

Work done = 900 MW

heat lost = 2100 MW

total energy from the hot reservior , Qh = 2100 + 900

Qh = 900 MW

Now , let the temperature of the combustion of the location is Th

W/Qh = 1 -Tc/Th

900/3000 = 1 - (15 + 273)/Th

Th = 411.4 K = 138.3 degree C

the minimum temperature that the combustion locaiton could possibly have is 138.3 degree C